Swimming space...
The length of a rectansular swimming pool is 2x - 1 meters, and thewidth is x + 2 meters. Write a polynomial A(x) that represents the area. Find A(5).

Answer by Alan3354(69443)

Swimming space...
The length of a rectansular swimming pool is 2x - 1 meters, and thewidth is x + 2 meters. Write a polynomial A(x) that represents the area.
Area = Lenght * width
A(x) = (2x-1)*(x+2) = 2x^2 + 3x - 2 sq meters
-------------
Find A(5)
Sub 5 for x:
A(5) = 2*5^2 + 3*5 - 2
A(5) = 50 + 15 - 2 = 63 sq meters
---------------
Note that you can also sub 5 in the Length and width:
A(5) = (2*5 - 1)*(5 + 2) = 9*7 = 63 sq meters